OpenStudy (anonymous):

What value of k makes the equation true? k - 26.7 = 12.8 A. 13.9 B. 38.5 C. 39.5 D. 39.9 @sammixboo

OpenStudy (misty1212):

HI!

OpenStudy (anonymous):

hey

OpenStudy (misty1212):

\[k – 26.7 = 12.8\\ k=12.8+26.7\] let me get a calculator

OpenStudy (misty1212):

\[12.8+26.7=39.5\]

sammixboo (sammixboo):

It is saying a number minus 26.7 is 12.8, and to find this number, you must add 26.7 and 12.8

OpenStudy (anonymous):

IS IT c.???

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]yes it is